OpenStudy (anonymous):

verify [cos(2x) -sin(2x)][sin(2x) + cos(2x)] = cos(4x)

OpenStudy (anonymous):

clue: (a-b)(a+b)=a^2-b^2 cos^2x-sin^2x=cos2x